# EditBoard: Towards A Comprehensive Evaluation Benchmark for Text-based Video Editing Models

### v1

@@ Abstract
The rapid development of diffusion models has significantly advanced AI-generated content (AIGC), particularly in Text-to-Image (T2I) and Text-to-Video (T2V) generation. Text-based video editing, leveraging these generative capabilities, has emerged as a promising field, enabling precise modifications to videos based on text prompts. Despite the proliferation of innovative video editing models, there is a conspicuous lack of comprehensive evaluation benchmarks that holistically assess these models’ performance across various dimensions. Existing evaluations are limited and inconsistent, typically summarizing overall performance with a single score, which obscures models’ effectiveness on individual editing tasks. To address this gap, we propose EditBoard, the first comprehensive evaluation benchmark for text-based video editing models. EditBoard encompasses nine automatic metrics across four dimensions, evaluating models on four task categories and introducing three new metrics to assess fidelity. This task-oriented benchmark facilitates objective evaluation by detailing model performance and providing insights into each model’s strengths and weaknesses. By open-sourcing EditBoard, we aim to standardize evaluation and advance the development of robust video editing models.
@@ 4.1 Data Preparation
We utilize samples from the DAVIS dataset (Pont-Tuset et al. 2017) to obtain the masks required for conducting Semantic Score testing. We also select samples from the LOVEU-TGVE-2023 dataset (Wu et al. 2023b). For each task, we select 10 videos containing a variety of objects, such as cars, animals, and humans. Each original video is paired with at least two target prompts according to the task category. For generating source prompts, we employ BLIP-2 (Li et al. 2023b) for the automated generation of video captions. The original frames are resized to a uniform resolution of 512$\times$512 to match the configuration of the testing models. We also ensure that sufficient original videos meet the requirements for applying FF-$\alpha$, allowing for the full adoption of both FF-$\alpha$ and FF-$\beta$. Additionally, we adjust the target prompts for Single Object Single Attribute (SOSA) so that more than half of the edits focus on the foreground object to facilitate Semantic Score evaluation.
@@ 4.3 Human Alignment Experiment
We conduct a human alignment experiment to verify whether the automatic metrics align with human perception. We calculate the percentage of questions where human comparisons match the implications of the metrics, with the results presented in Table 2 (see the Appendix). Additionally, we assess whether the newly proposed FF-$\alpha$ and FF-$\beta$ metrics can effectively capture temporal flickering and motion consistency, assuming the original video has high quality. Therefore, edited videos with higher fidelity to the original should exhibit fewer flickering or inconsistency issues. The results are promising, indicating that videos with a higher FF score (i.e., worse quality) are consistently annotated as having more pronounced temporal flickering and motion inconsistency. Consequently, traditional metrics for generative models, such as motion smoothness, may not be necessary for video editing evaluation. Further demonstrations of the new metrics are presented in Appendix Section 8.
@@ 5.2 Trade-off Between Execution and Fidelity
5.3 Style Editing Helps Execution
An interesting observation from our experiments is that some models tend to achieve higher execution scores on SOMA tasks than SOSA tasks, despite the former’s increased complexity. This anomaly is partly due to some SOMA target prompts involving editing multiple attributes along with the global style. Upon reviewing the results, we find that models execute object editing more effectively when the style becomes more abstract. For example, FateZero struggles to transform a man into an ape in SOMA. However, when an additional prompt to apply a Van Gogh style is added, the execution improves significantly (see Figure 5). Future research can leverage this finding to explore how style editing can enhance execution, potentially offering new avenues for improving model performance. More discussions are provided in Appendix Section 7.
Figure 5: Despite the structural change in Edit Result 2, the model successfully turns the man into an ape with the additional prompt of applying Van Gogh painting style.
@@ 6 Conclusion
In this paper, we propose EditBoard, a pioneering comprehensive evaluation benchmark specifically designed for text-based video editing models. Our benchmark addresses the critical need for a standardized framework that holistically assesses the multifaceted performance of these models. By incorporating nine metrics across four dimensions and introducing three novel metrics, EditBoard provides a detailed, task-oriented evaluation that highlights each model’s strengths and weaknesses. Empirical results demonstrate EditBoard’s efficacy in aligning with human perceptions of video quality and editing precision. By open-sourcing EditBoard, we aim to foster the development of more robust and reliable video editing models, ultimately advancing the evolving field of AIGC. Our work sets a new standard for evaluating text-based video editing models, ensuring a more comprehensive and objective assessment for future research.
Appendix
7 Additional Evaluation Results
7.1 Ground-A-Video’s Accurate Editing
From the results presented in Table 1, Ground-A-Video, a model designed for multi-attribute editing tasks, demonstrates high accuracy in identifying the target object for editing, as evidenced by its Semantic Score. Additionally, the model achieves high Success Rate and CLIP Similarity scores, indicating effective execution capabilities. The lower FF-$\alpha$ and FF-$\beta$ scores further suggest strong fidelity. This robust performance can be largely attributed to the model’s integration of grounding information from GLIP (Li et al. 2022), which provides precise localization of the object to be edited. Additionally, incorporating inflated ControlNet (Zhang, Rao, and Agrawala 2023) enhances frame-to-frame consistency and structural fidelity to the original video. Despite the extra work requiring users to acquire grounding information and feed it to the model, Ground-A-Video delivers superior editing performance.
7.2 Results of TokenFlow and Video-P2P
With the tuning stage, Video-P2P performs better at object-aware editing, compared to FateZero which is a zero-shot model but has an attention leakage problem. However, Video-P2P struggles with SOMA and MOA tasks, showing limitations in multi-attribute editing. By enforcing consistency within the diffusion feature space, TokenFlow performs well in the dimension of fidelity, indicating that the edited videos retain their original structure and motion. Yet, its execution remains deficient, as it occasionally fails to achieve multi-attribute editing.
7.3 Per-task Results
We visualize the per-task performance of the five models in Figure 10 and Tables 3456. Except for the overall strong performance of Ground-A-Video, TokenFlow performs better on SOSA and MOA tasks. Both FateZero and Control-A-Video demonstrate commendable performance in SE and SOMA tasks. Moreover, CLIP Similarity exhibits limited differentiation capability, leading to similar scores across all tested models. This issue could potentially be attributed to the visual limitations of CLIP. CLIP primarily focuses on global semantics and struggles to distinguish fine visual differences in similar images (Tong et al. 2024). Since the original frame largely dictates the global structure of the edited frame, some editing results appear too similar for CLIP to differentiate effectively.
Figure 6: A Sample Question for Human Alignment Experiment
8 More Results of Newly Proposed Metrics
We visualize pairs of edited frame sequences generated by different models using the same source prompt and original video. The questions used in the human alignment experiment are similar to the presented comparison, with one sample question shown in Figure 6. The newly proposed FF-$\alpha$ and FF-$\beta$ metrics effectively reflect the fidelity of edited frames to the original frames. For instance, in the bear sample illustrated in Figure 7, Edit Result 1 shows a noticeable difference in the position of the bear’s head, which is detected by FF-$\alpha$, leading to a higher (worse) score. Similarly, Edit Result 2 of the airplane sample exhibits frame-wise background inconsistency, indicated by the red bounding box, which results in a higher FF-$\alpha$ score. For FF-$\beta$, Edit Result 1 of the car-turn sample in Figure 8 reveals an inconsistent shape of the car, reflected in a higher FF-$\beta$ score. Additionally, the subtle misalignment in Edit Result 2 of the swan sample is also captured by a higher FF-$\beta$ score.
Moreover, we present visualizations for Semantic Score comparisons, demonstrating that Semantic Score is an exemplary metric to gauge a model’s precision in editing. As shown in Figure 9, Edit Result 2 of the car-turn sample shows that the road suffers from minor changes to red, which results in a higher semantic score. Similarly, the subtle difference in Edit Result 2 of the school bus sample is also reflected by a higher Semantic Score. In summary, the newly proposed metrics align well with human perception and contribute to a more comprehensive and unified evaluation benchmark.
9 Limitation and Future Work
While our task-oriented evaluation benchmark contributes to the standardized and unified assessment of text-based video editing models, it has certain limitations. As video editing tasks continue to evolve alongside advancements in video editing models, new tasks such as object deletion from videos have emerged. We plan to continuously incorporate additional tasks to ensure a comprehensive evaluation. Furthermore, with the development of Vision-Language Learning Models (VLLMs) such as Otter (Li et al. 2023a), BLIPv2 (Li et al. 2023b), and EMU (Sun et al. 2023), there is potential to leverage their vision question answering and video captioning capabilities to enhance frame-text alignment evaluation. Exploring these possibilities will be a key focus for future research.
Figure 7: Visual comparison of different edited frames’ FF-$\alpha$ score.
Figure 8: Visual comparison of different edited frames’ FF-$\beta$ score.
Figure 9: Visual comparison of different edited frames’ Semantic Score.
  FF-$\alpha$ FF-$\beta$CenterstackSemantic
[…]
Similarity
Matching Rate 92.54 89.93 95.24 92.68 85.72
Table 2: Percentage of questions that humans give the same result as automatic metrics. It shows that the automatic metrics generally align with human perception.
Figure 10: Visualization of the five model’s performance over each task. SOSA: Single Object Single Attribute; SE: Style Editing; SOMA: Single Object Multiple Attributes; MOA: Multiple Objects and Attributes.
[…]
Quality $\uparrow$
  FateZero 8.0082 0.1723 25.2665 0.5294 0.3105 0.9696 0.9497 0.5546 0.6907
Control-A-Video 18.0534 0.2674 66.1538 0.6050 0.3170 0.9672 0.9700 0.5377 0.6973
Ground-A-Video 6.0249 0.1022 8.2680 0.8659 0.3239 0.9622 0.9711 0.5635 0.6750
Video-P2P 11.8893 0.2216 20.7714 0.5156 0.3037 0.9692 0.9696 0.4847 0.6665
TokenFlow 7.2708 0.1566 31.6023 0.6471 0.3242 0.9790 0.9525 0.6233 0.7408
Table 3: Model performance on the task of SOSA.
[…]
Quality $\uparrow$
  FateZero 10.5420 0.5022 - 0.6923 0.3341 0.9417 0.9563 0.5886 0.6267
Control-A-Video 11.9252 0.5993 - 0.7143 0.3212 0.9440 0.9626 0.5575 0.7087
Ground-A-Video 7.7620 0.4259 - 0.9443 0.3452 0.9706 0.9503 0.5703 0.6914
Video-P2P 12.0104 0.5463 - 0.6058 0.3113 0.9561 0.9704 0.5125 0.5962
TokenFlow 6.3735 0.4364 - 0.4135 0.3123 0.9762 0.9629 0.6421 0.7010
Table 4: Model performance on the task of SE.
[…]
Quality $\uparrow$
  FateZero 10.2624 0.2795 - 0.6667 0.3108 0.9418 0.9280 0.5086 0.6083
Control-A-Video 15.3962 0.3977 - 0.8429 0.3251 0.9595 0.9597 0.5751 0.7006
Ground-A-Video 7.3247 0.2293 - 0.8723 0.3479 0.9313 0.9486 0.5676 0.6673
Video-P2P 12.3277 0.3441 - 0.3462 0.2764 0.9657 0.9546 0.4827 0.6315
TokenFlow 7.8735 0.2437 - 0.5750 0.3125 0.9739 0.9437 0.5842 0.6973
Table 5: Model performance on the task of SOMA.
[…]
Quality $\uparrow$
  FateZero 12.2363 0.2832 - 0.4464 0.2944 0.9457 0.9449 0.4970 0.4867
Control-A-Video 21.2607 0.5616 - 0.7347 0.3025 0.9249 0.9485 0.5085 0.7137
Ground-A-Video 7.2733 0.2337 - 0.8370 0.3340 0.9687 0.9502 0.5681 0.6737
Video-P2P 9.1412 0.4087 - 0.3752 0.2925 0.9533 0.9606 0.4921 0.5383
TokenFlow 8.3205 0.3098 - 0.5532 0.3084 0.9546 0.9595 0.5535 0.6759
Table 6: Model performance on the task of MOA.

### v2

@@ Abstract
The rapid development of diffusion models has significantly advanced AI-generated content (AIGC), particularly in Text-to-Image (T2I) and Text-to-Video (T2V) generation. Text-based video editing, leveraging these generative capabilities, has emerged as a promising field, enabling precise modifications to video content based on textual prompts. Despite the proliferation of innovative video editing models, there is a conspicuous lack of comprehensive evaluation frameworks that holistically assess these models’ performance across various dimensions. Existing metrics are limited, inconsistent, and focused on assigning a single score per metric, failing to reveal model’s performance on each editing task. To address this gap, we propose EditBoard, the first comprehensive evaluation benchmark for text-based video editing models. EditBoard encompasses nine automatic metrics across four key dimensions, evaluating models on four categories of tasks, and introduces three new metrics to assess fidelity. This task-oriented framework facilitates objective evaluation by breaking down model performance into details, providing insights into each model’s strengths and weaknesses. By open-sourcing EditBoard, we aim to standardize evaluation and advance the development of robust video editing models.
Codehttps://github.com/Samchen2003/EditBoard
@@ 4.1 Data Preparation
We utilize samples from the DAVIS dataset (Pont-Tuset et al. 2017) to obtain the masks required for conducting Semantic Score testing. We also select samples from the LOVEU-TGVE-2023 dataset (Wu et al. 2023b). For each task, we select 10 videos containing a variety of objects, such as cars, animals, and humans. Each original video is paired with at least two target prompts according to the task category. For generating source prompts, we employ BLIP-2 (Li et al. 2023) for the automated generation of video captions. The original frames are resized to a uniform resolution of 512$\times$512 to match the configuration of the testing models. We also ensure that sufficient original videos meet the requirements for applying FF-$\alpha$, allowing for the full adoption of both FF-$\alpha$ and FF-$\beta$. Additionally, we adjust the target prompts for Single Object Single Attribute (SOSA) so that more than half of the edits focus on the foreground object to facilitate Semantic Score evaluation.
@@ 4.3 Human Alignment Experiment
We conduct a human alignment experiment to evaluate whether the automatic metrics align with human perception. This is done by calculating the percentage of questions where human judgments match the outcomes suggested by the metrics, as shown in Table 2. In this experiment, human annotators compare two samples based on a specific evaluation criterion and select the one with higher quality or indicate that the difference is negligible. A match is recorded if the human choice agrees with the metrics’ comparison. Also, if annotators mark the samples as “hard to distinguish” and the metrics’ values differ by less than a predefined threshold, it is also considered a match.
  FF-$\alpha$ FF-$\beta$CenterstackSemantic
[…]
Similarity
Matching Rate 92.54 89.93 95.24 92.68 85.72
Table 2: Our new metrics achieve around 90% match with human choices.
We also provide the Pearson’s correlation score in Table 3 to further validate that our new metrics align with human perception.
  FF-$\alpha$ FF-$\beta$CenterstackSemantic
[…]
Similarity
Pearson’s score 0.9312 0.9126 0.9417 0.9333 0.8769
Table 3: Pearson correlation scores demonstrate a similar trend, reinforcing the alignment.
Additionally, we evaluate FF-$\alpha$ and FF-$\beta$ for capturing temporal flickering and motion consistency, finding that higher FF scores (lower quality) align with more pronounced issues. This suggests traditional metrics like motion smoothness may be unnecessary for video editing evaluation.
@@ 6 Conclusion
In this paper, we propose EditBoard, a pioneering comprehensive evaluation benchmark specifically designed for text-based video editing models. Our benchmark addresses the critical need for a standardized framework that holistically assesses the multifaceted performance of these models. By incorporating nine metrics across four dimensions and introducing three novel metrics, EditBoard provides a detailed, task-oriented evaluation that highlights each model’s strengths and weaknesses. Empirical results demonstrate EditBoard’s efficacy in aligning with human perceptions of video quality and editing precision. By open-sourcing EditBoard, we aim to foster the development of more robust and reliable video editing models, ultimately advancing the field of AIGC. Our work sets a new standard for evaluating text-based video editing models, ensuring a more comprehensive and objective assessment for future research.
Acknowledgments
We sincerely thank the reviewers for their constructive feedback and insightful suggestions, which have greatly improved this work. We also extend our heartfelt gratitude to Prof. Baoyuan Wu, Prof. Mingqiang Wei, Prof. Limin Wang, and Dr. Peng Zhou for their invaluable guidance and support throughout this research.